OpenStudy (anonymous):

Solve.x+2 over 15 + 3 minus x over 5 = 2 over 3 A. x = 1 B. x = 0.5 C. x = 2

OpenStudy (radar):

It's B.

OpenStudy (radar):

B checks out correctly. (x+2+9-3x)/15 = 2/3 (-2x+11)/15 = 2/3 -6x+33=30 -6x=-3 x=1/2 or 0.5